# HyperspectralViTs: Fast and Accurate methane detection on-board satellites

### v1

@@ Abstract
On-board processing of hyperspectral data with machine learning models would enable unprecedented amount of autonomy for a wide range of tasks, for example methane detection or mineral identification. Methane is the second most important greenhouse gas contributor to climate change, and it’s automated detection on-board of satellites using machine learning models would allow for early warning system and could enable new capabilities such as automated scheduling inside constellations of satellites. Classical methods for methane detection suffer from high false positive rates and previous deep learning models exhibit prohibitive computational requirements. We propose fast and accurate machine learning architectures which support end-to-end training with data of high spectral dimension. We evaluate our models on two tasks related to hyperspectral data processing - methane leak detection and mineral identification. With our proposed general architectures, we improve the F1 score of the previous methane detection state-of-the-art models by more than 27% on a newly created synthetic dataset and by almost 13% on the previously released large benchmark dataset. We also demonstrate that training models on the synthetic dataset improves performance of models finetuned on the dataset of real events by 6.9% in F1 score in contrast with training from scratch. On a newly created dataset for mineral identification, our models provide 3.5% improvement in the F1 score in contrast to the default versions of the models. With our proposed models we improve the inference speed by 85.19% in contrast to previous classical and deep learning approaches by removing the dependency on classically computed features. Namely, one capture from the EMIT sensor can be processed in only 30 seconds on a realistic proxy hardware used on the ION-SCV 004 satellite.
@@ I Introduction
Mineral identification is a task related to scene decomposition into individual endmembers from a spectral library. On its own it could serve as a strong general tasks highly relevant for future Foundation models working with hyperspectral data. We note that one of the example downstream tasks of interest could also be methane detection, as it has been shown that having information about scene background is beneficial for separation of methane leak events from confounders [4].
@@ I-A Literature review
As has been identified by [29] and several overview papers [30, 31, 32], research in machine learning for processing hyperspectral data has been limited mainly due to the lack of large, publicly available datasets with high quality annotations. Many works still use very small datasets such as Indian Pines [33], University of Pavia, Houston or Salinas, all of which contain only a few data captures. As further noted by [30], this data is often divided into almost equal train and test subsets, which leads to model overfitting and results with very high accuracies almost regardless of the used model. In the domain of methane leak detection, only the recent works of [34, 4] release processed datasets from raw data captures, [34] furthermore refines the available labels. The works of [35, 36] propose machine learning models for detection and quantification of methane leak events in hyperspectral data, however the used datasets and codes aren’t published. In contrast, we publish all our used code and trained models and present the created datasets in a machine learning ready format.
@@
Deployment of machine leaning models on-board of satellites with hyperspectral sensors was shown in [39] to filter cloudy data captures. The demonstrator Earth Observer 1 (EO-1) mission by NASA hosted simple thresholding models that processed hyperspectral data on-board to detect clouds and pre-classicy scenes [40] and later used models for mineral identification in [41]. On-board deployment of unsupervised models to detect anomalous events in multispectral data was proposed in [42] and later tested on-board of a real satellite in [43]. These examples motivate us that methane detection using deep learning models on the current generation of hardware on today’s satellites would be feasible. They also highlight the potential benefits from autonomy on-board.
@@ II-A Methane leak event simulation
In one of our datasets, we simulate synthetic methane leak events in the clean hyperspectral datacubes, given the knowledge of the expected signal of methane and plume concentration data from real events. Unlike previous study of [44], which used plume shapes simulated using the expansive Large Eddy Simulation (LES) [45], we use real methane leak events from the similar hyperspectral sensor AVIRIS-NG, to inform the simulation. Namely, we use the STARCOP dataset [34], given it’s improved and manually refined labels and high diversity of plume sizes.
@@ II-B Mineral aggregation
To create our new mineral dataset, we adapt the steps described in the Theoretical basis document [46] released alongside the “EMITL2BMIN.001” product [47] computed from the raw EMIT data. This data contains identifications of 381 individual components (separated into two groups). These can be aggreggated into larger mineral classes, for this we select the so-called “EMIT10” subset defined in [46]. The aggregation process is shown on Figure 4. We note that we use only the binary maps of three selected mineral classes out of the total 10. We chose the subset M = {Goethite, Hematite, Kaolinite}, as these classes were abundant in the locations sampled in our dataset. We use the released annotations which were computed using the Tetracorder method [24], and as such we can’t really use Tetracorder as a classical baseline (as it would unfairly obtain perfect scores). Therefore, we consider the created dataset as having pseudo-ground truth labels, and our results only as initial. We however note, that given any manually validated dataset, our models could easily be retrained with the new data.
@@ II-C Final created datasets
Firstly, we created the “OxHyperSyntheticCH4” dataset of synthetically simulated methane leak events using the labels of the STARCOP dataset with newly downloaded near-global hyperspectral datacubes from the EMIT sensor. In total we keep 796 large tiles (of 512x512 px) in the training dataset, 198 in the validation dataset and 200 in the test dataset. Each subset is made from completely non-overlapping sources of data (clean EMIT datacubes and labels from the STARCOP dataset from AVIRIS-NG data). In each subset we uniformly sample the available plume sizes, so that we have similar distribution of event sizes. Half of these tiles contain simulated events while the other half is kept event free to provide negative samples. We note that these large tiles (of 512x512 px) are further tiled when training the models. When using the tile size of 64 pixels with an overlap of 32 pixels, we finally get a training dataset of 174341 samples. The total size of this data is 228 GB.
Secondly, we create the “OxHyperRealCH4” dataset of real methane leak events. In total we have 279 large tiles in the training dataset, 91 in the validation dataset and 98 in the test dataset. These have reliable ground truth labels, which we calculated from the original labels provided on the EMIT data portal and consequentially manually checked. We note that this dataset is smaller than the other ones and training a machine learning model from scratch could lead to overfitting and poor generalisation, however it can also be used for model fine-tuning and to test generalisation ability of models trained on other datasets. The total size of this data is 47 GB.
Thirdly, we create “OxHyperMinerals”, a completely new dataset for mineral identification, with labels of 3 selected mineral classes and source annotation for all 381 constituents (separated into two groups). In total we have 796 large tiles in the training dataset, 198 in the validation and finally 200 in the test dataset. These large tiles exactly match the tiles used in the synthetic methane events dataset, however they do not contain the simulated events and contain more bands. The mineral dataset contains all 285 spectral bands in between spectral range of 381-2492nm. The total size of this data is 372 GB.
@@ III-A Limitations of existing approaches
We are using three different semantic segmentation architectures with the created dataset of hyperspectral datacubes - the commonly used U-Net model [48] and the newer SegFormer [1] and EfficientViT [2] models. These models can be understood in terms of the encoder-decoder layout. The U-Net model consists of downsampling convolutional layers in the encoder network, and upsampling layers in the decoder, while also using skip-connections to communicate across the same spatial resolution of the data. The SegFormer model uses Transformer blocks in its encoder network, concatenates the features from different resolution levels and then uses a MLP block in the decoder network. The EfficientViT models further proposes changes to make the model run faster on limited hardware. In all models the spatial resolution is initially reduced and then reconstructed in the decoder, while information at different scales is kept with the use of skip-connections, or when concatenating the features together.
@@ IV Experimental Setup
Finally, when training models on the “OxHyperMinerals” dataset on the task of mineral detection, we train all models for 20 epochs with tiles of 64 px with 32 px overlap. In this case, trained models use all 285 bands of the available EMIT data (381-2492 nm). In this case, the data loader is not balancing the samples and losses are not weighted.
@@ IV-A Hardware Restrictions
Therefore, with several tested model variants, we also conducted experiments measuring the runtime on compute constrained hardware, namely on the flatbed hardware provided by Unibap and D-Orbit [37]. These machines serve as a proxy to match the computational in-space environment of a D-Orbit’s ION-SCV004 satellite, which is equipped with: quad-core 1.2GHz, AMD GX-412HC SOC CPU processor, the Intel Movidius Myriad X VPU and 2GB of RAM. Similar compute environments leveraging Myriad chips were used on prior missions such as ESA’s PhiSat-1 [51] and PhiSat-2 [52].
@@ V-B Methane detection on AVIRIS-NG
We have encountered very large gains on the synthetic and real EMIT datasets, however smaller ones on the benchmark dataset of AVIRIS-NG data. We note that perhaps better versions of the MF product could lead to stronger baselines also in the EMIT data. For example [57] explores scene specific adaptation of MF products, and similarly there could be some EMIT specific improvements for calculating MF. This difference could also be caused by the number of used bands - while on the STARCOP dataset the end-to-end models use 60 bands, on the new EMIT datasets we use 86 bands which also demonstrate a wider spectral range coverage.
@@ V-D Inference times on low compute hardware
TABLE V: Timing of models on different devices. We estimate the time in seconds for processing a full granule of EMIT data (1280x1242px tiled into 100 tiles of 128x128px). EfficientViT models weren’t evaluated on the Unibap machine. IO time to load the hyperspectral tiles is kept separate in the last row.
Unibap CPU Rasp3 CPU Rasp4 CPU Xavier CPU X. GPU (onnx) X. GPU (trtexec)
HyperSTARCOP MF+RGB [34] 203.3 1025.25 286.57 377 371.17 370.27
SegFormer Base 15.3 38.82 9.19 7.87 1.62 0.49
SegFormer UpConv 30.1 105.43 28.7 13.92 1.97 0.64
SegFormer UpConvStride 119.5 399.76 107.87 48.38 2.95 1.37
EfficientViT Base N/A 39.96 9.45 9.39 2.36 0.37
EfficientViT ConvUp N/A 47.73 13.81 9.99 2.00 0.55
EfficientViT ConvUpStride N/A 91.07 21.26 18.09 2.10 0.5
IO 4 39.99 12.88 6.13 6.13 6.13
@@ VI Conclusion
To conclude, we demonstrate new state-of-the-art results in methane detection using our end-to-end machine learning architectures adapted for hyperspectral data. Namely, on a newly created dataset of synthetic events in data from the EMIT sensor, we improve the F1 score by more than 27% against the prior models depending on classical methane enhancement products from [34] and by over 97% in F1 score over classical matched filter baselines. When comparing with off-the-shelf performance, our proposed Hyperspectral variant of the SegFormer model improves the F1 score by over 23% on this synthetic dataset. We confirm our findings also on a benchmark dataset released by [34], where we improve the overall F1 score by almost 13% and the AUPRC score by more than 17%. Finally, on a newly created dataset of real methane leak events recorded by the EMIT sensor, we demonstrate, that it is possible to leverage models pretrained on synthetic data by fine-tuning them on smaller real datasets. Using the best performing model variant, we show that the finetuned model improves the F1 score by 6.9% over models trained from scratch.
@@
As a potential limitation of the proposed end-to-end Hyper SegFormer model, it operates directly on the raw hyperspectral data and as such, it can’t be used for zero-shot generalisation as for example the model of [34]. We however note, that a similar simulated dataset can be created from representative data for any other satellite and later used to train end-to-end models. Secondly, our model doesn’t support additional tasks such as plume quantification or source point estimation. We however note, that the prediction of our model can be used as a binary mask to clean the noisy methane enhancement products, after which classical pipelines for these tasks can be used [62].
For future research, we’d like to explore further steps to make the models more efficient. Our proposed changes to the Transformer based models model can be combined with pruning reported in [63, 38]. Secondly, we’d also like to explore extension of our models with methods for uncertainty estimation. Methods such as model ensembles [64], or Monte Carlo Dropout [65] could be easily added to the proposed semantic segmentation models, as was previously demonstrated in [66]. Thirdly, we note that the existing archives of EMIT data (currently with more than 107k captures totaling more than 215 TB of data) offer unique opportunities to train large general foundation models on hyperspectral data, where one of the example tested tasks could also be methane detection. Models proposed in this paper could likely be leveraged for general learning tasks (in the direction outlined by our mineral identification task), but also for other downstream applications highlighted by the future planned missions such as NASA’s SBG [58] or ESA’s CHIME [67].
Finally, we are releasing the three newly created datasets on [TBR], [TBR] and [TBR] and the code and the pre-trained models on Github at [TBR].

### v2

# HyperspectralViTs: General Hyperspectral Models for On-board Remote Sensing

@@ Abstract
On-board processing of hyperspectral data with machine learning models would enable unprecedented amount of autonomy for a wide range of tasks, for example methane detection or mineral identification. This can enable early warning system and could allow new capabilities such as automated scheduling across constellations of satellites. Classical methods suffer from high false positive rates and previous deep learning models exhibit prohibitive computational requirements. We propose fast and accurate machine learning architectures which support end-to-end training with data of high spectral dimension without relying on hand-crafted products or spectral band compression preprocessing. We evaluate our models on two tasks related to hyperspectral data processing. With our proposed general architectures, we improve the F1 score of the previous methane detection state-of-the-art models by 27% on a newly created synthetic dataset and by 13% on the previously released large benchmark dataset. We also demonstrate that training models on the synthetic dataset improves performance of models finetuned on the dataset of real events by 6.9% in F1 score in contrast with training from scratch. On a newly created dataset for mineral identification, our models provide 3.5% improvement in the F1 score in contrast to the default versions of the models. With our proposed models we improve the inference speed by 85% in contrast to previous classical and deep learning approaches by removing the dependency on classically computed features. With our architecture, one capture from the EMIT sensor can be processed within 30 seconds on realistic proxy of the ION-SCV 004 satellite.
@@ I Introduction
Mineral identification is a task related to scene decomposition into individual endmembers from a spectral library. On its own it could serve as a strong general task highly relevant for future Foundation models working with hyperspectral data. We note that one of the example downstream tasks of interest could also be methane detection, as it has been shown that having information about scene background is beneficial for separation of methane leak events from confounders [4].
@@ I-A Literature review
As has been identified by [29] and several overview papers [30, 31, 32], research in machine learning for processing hyperspectral data has been limited mainly due to the lack of large, publicly available datasets with high quality annotations. Many works still use very small datasets such as Indian Pines [33], University of Pavia, Houston or Salinas, all of which contain only a few data captures. As further noted by [30], this data is often divided into almost equal train and test subsets, which leads to model overfitting and results with very high accuracies almost regardless of the used model. In the domain of methane leak detection, only the recent works of [34, 4] release processed datasets from raw data captures, [34] furthermore refines the available labels. The works of [35, 36] propose machine learning models for detection and quantification of methane leak events in hyperspectral data, however the used datasets and code/model weights are not published. In contrast, we publish all our used code and trained models and present the created datasets in a machine learning ready format.
@@
Deployment of machine leaning models on-board of satellites with hyperspectral sensors was shown in [39] to filter cloudy data captures. The demonstrator Earth Observer 1 (EO-1) mission by NASA hosted simple thresholding models that processed hyperspectral data on-board to detect clouds and pre-classify scenes [40] and later used models for mineral identification in [41]. On-board deployment of unsupervised models to detect anomalous events in multispectral data was proposed in [42] and later tested on-board of a real satellite in [43]. These examples motivate us that methane detection using deep learning models on the current generation of hardware on today’s satellites would be feasible. They also highlight the potential benefits from autonomy on-board in future satellite constellations.
@@ II-A Methane leak event simulation
In one of our datasets, we simulate synthetic methane leak events in the clean hyperspectral datacubes, given the knowledge of the expected signal of methane and plume concentration data from real events. Unlike the previous study of [44], which used plume shapes simulated using the expansive Large Eddy Simulation (LES) [45], we use real methane leak events from the similar hyperspectral sensor AVIRIS-NG, to inform the simulation. Namely, we use the STARCOP dataset [34], given it’s improved and manually refined labels and high diversity of plume sizes.
@@ II-B Mineral aggregation
To create our new mineral dataset, we adapt the steps described in the Theoretical basis document [46] released alongside the “EMITL2BMIN.001” product [47] computed from the raw EMIT data. This data contains identifications of 381 individual components (separated into two groups). These can be aggreggated into larger mineral classes, for this we select the so-called “EMIT10” subset defined in [46]. The aggregation process is shown on Figure 4. We note that we use only the binary maps of three selected mineral classes out of the total 10. We chose the subset M = {Goethite, Hematite, Kaolinite}, as these classes were abundant in the locations sampled in our dataset. We use the released annotations which were computed using the Tetracorder method [24], and as such we can’t really use Tetracorder as a classical baseline (as it would unfairly obtain perfect scores). Therefore, we consider the created dataset as having pseudo-ground truth labels, and our results only as initial proof-of-principle. We however note, that given any manually validated dataset, our models could be easily retrained with the new data.
@@ II-C Final created datasets
Firstly, we created the “OxHyperSyntheticCH4” dataset of synthetically simulated methane leak events using the labels of the STARCOP dataset with newly downloaded near-global hyperspectral datacubes from the EMIT sensor (in L1B level of data processing). In total we keep 796 large tiles (of 512x512 px) in the training dataset, 198 in the validation dataset and 200 in the test dataset. Each subset is made from completely non-overlapping sources of data (clean EMIT datacubes and labels from the STARCOP dataset from AVIRIS-NG data). In each subset we uniformly sample the available plume sizes, so that we have similar distribution of event sizes. Half of these tiles contain simulated events while the other half is kept event free to provide negative samples. We note that these large tiles (of 512x512 px) are further tiled when training the models. When using the tile size of 64 pixels with an overlap of 32 pixels, we finally get a training dataset of 174341 samples. The total size of this data is 228 GB.
Secondly, we create the “OxHyperRealCH4” dataset of real methane leak events. In total we have 279 large tiles in the training dataset, 91 in the validation dataset and 98 in the test dataset. When further tiled, the training dataset has 54805 samples (64 pixel tiles). These have reliable ground truth labels, which we calculated from the original labels provided on the EMIT data portal and consequentially manually checked. We note that this dataset is smaller than the other ones and training a machine learning model from scratch could lead to overfitting and poor generalisation, however it can also be used for model fine-tuning and to test generalisation ability of models trained on other datasets. The total size of this data is 47 GB.
Thirdly, we create “OxHyperMinerals”, a completely new dataset for mineral identification, with labels of 3 selected mineral classes and source annotation for all 381 constituents (separated into two groups). In total we have 796 large tiles in the training dataset, 198 in the validation and finally 200 in the test dataset. When further tiled, the training dataset has 174341 samples (64 pixel tiles). The large tiles exactly match the tiles used in the synthetic methane events dataset, however they do not contain the simulated events and contain more bands. The mineral dataset contains all 285 spectral bands in between spectral range of 381-2492nm. The total size of this data is 372 GB.
@@ III-A Limitations of existing approaches
We are using three different semantic segmentation architectures with the created dataset of hyperspectral datacubes - the commonly used U-Net model [48] and the newer SegFormer [1] and EfficientViT [2] models. These models can be understood in terms of the encoder-decoder layout. The U-Net model consists of downsampling convolutional layers in the encoder network, and upsampling layers in the decoder, while also using skip-connections to communicate across the same spatial resolution of the data. The SegFormer model uses Transformer blocks in its encoder network, concatenates the features from different resolution levels and then uses a MLP block in the decoder network. The EfficientViT model further proposes changes to make the model run faster on limited hardware. In all models the spatial resolution is initially reduced and then reconstructed in the decoder, while information at different scales is kept with the use of skip-connections, or when concatenating the features together.
@@ IV Experimental Setup
Finally, when training models on the “OxHyperMinerals” dataset on the task of mineral detection, we train all models for 20 epochs with tiles of 64 px with 32 px overlap. In this case, trained models use all 285 bands of the available EMIT data (381-2492 nm). Here, the data loader is not balancing the samples and losses are not weighted.
@@ IV-A Hardware Restrictions
Therefore, with several tested model variants, we also conducted experiments measuring the runtime on compute constrained hardware, namely on the flatbed hardware provided by Unibap and D-Orbit [37]. These machines serve as a proxy to match the computational in-space environment of a D-Orbit’s ION-SCV004 satellite, which is equipped with: quad-core 1.2GHz, AMD GX-412HC SOC CPU processor, the Intel Movidius Myriad X VPU and 2GB of RAM. The same compute environment was used in [43] and similar environments leveraging Myriad chips were used on prior missions such as ESA’s PhiSat-1 [51] and PhiSat-2 [52].
@@ V-B Methane detection on AVIRIS-NG
We have encountered very large gains on the synthetic and real EMIT datasets, however smaller ones on the benchmark dataset of AVIRIS-NG data. We note that perhaps better versions of the MF product could lead to stronger baselines also in the EMIT data. For example [57] explores scene specific adaptation of MF products, and similarly there could be some EMIT specific improvements for calculating MF. This difference could also be caused by the number of used bands - while on the STARCOP dataset the end-to-end models use 60 bands, on the new EMIT datasets we use 86 bands which also have a wider spectral range coverage.
@@ V-D Inference times on low compute hardware
TABLE V: Timing of models on different devices. We estimate the time in seconds for processing a full granule of EMIT data (1280x1242px tiled into 100 tiles of 128x128px). EfficientViT models weren’t evaluated on the Unibap machine. IO time to load the hyperspectral tiles is kept separate in the last row. We also illustrate the number of parameters each model uses.
Params. (M) Unibap CPU Rasp3 CPU Rasp4 CPU Xavier CPU X. GPU (onnx) X. GPU (trtexec)
HyperSTARCOP MF+RGB [34] 6.633 203.3 1025.25 286.57 377 371.17 370.27
SegFormer Base 3.845 15.3 38.82 9.19 7.87 1.62 0.49
SegFormer UpConv 4.326 30.1 105.43 28.7 13.92 1.97 0.64
SegFormer UpConvStride 4.326 119.5 399.76 107.87 48.38 2.95 1.37
EfficientViT Base 4.811 N/A 39.96 9.45 9.39 2.36 0.37
EfficientViT ConvUp 4.855 N/A 47.73 13.81 9.99 2.00 0.55
EfficientViT ConvUpStride 4.855 N/A 91.07 21.26 18.09 2.10 0.5
IO N/A 4 39.99 12.88 6.13 6.13 6.13
@@ VI Conclusion
Secondly, we demonstrate new state-of-the-art results in methane detection using our end-to-end machine learning architectures adapted for hyperspectral data. Namely, on a newly created dataset of synthetic events in data from the EMIT sensor, we improve the F1 score by more than 27% against the prior models depending on classical methane enhancement products from [34] and by over 97% in F1 score over classical matched filter baselines. When comparing with off-the-shelf performance, our proposed Hyperspectral variant of the SegFormer model improves the F1 score by over 23% on this synthetic dataset. We confirm our findings also on a benchmark dataset released by [34], where we improve the overall F1 score by almost 13% and the AUPRC score by more than 17%. Finally, on a newly created dataset of real methane leak events recorded by the EMIT sensor, we demonstrate, that it is possible to leverage models pretrained on synthetic data by fine-tuning them on smaller real datasets. Using the best performing model variant, we show that the finetuned model improves the F1 score by 6.9% over models trained from scratch.
@@
As a potential limitation of the proposed end-to-end HyperspectralViT models, they operate directly on the raw hyperspectral data and as such, they can’t be used for zero-shot generalisation as for example the model of [34]. We however note, that a similar simulated dataset can be created from representative data for any other satellite and later used to train end-to-end models. Secondly, our model doesn’t support additional tasks such as plume quantification or source point estimation. We however note, that the prediction of our model can be used as a binary mask to clean the noisy methane enhancement products, after which classical pipelines for these tasks can be used [62].
For future research, we’d like to explore further steps to make the models more efficient. Our proposed changes to the Transformer based models model can be combined with pruning reported in [63, 38]. Specialised accelerators as those suggested by [64] could also be leveraged. Secondly, we’d also like to explore extension of our models with methods for uncertainty estimation. Methods such as model ensembles [65], or Monte Carlo Dropout [66] could be easily added to the proposed semantic segmentation models, as was previously demonstrated in [67]. Thirdly, we note that the existing archives of EMIT data (currently with more than 107k captures totaling more than 215 TB of data) offer unique opportunities to train large general foundation models on hyperspectral data, where one of the example tested tasks could also be methane detection. Models proposed in this paper could likely be leveraged for general learning tasks (in the direction outlined by our mineral identification task), but also for other downstream applications highlighted by the future planned missions such as NASA’s SBG [58] or ESA’s CHIME [68].
Finally, we are releasing the three newly created datasets on https://huggingface.co/previtus [exact links TBR] and the code and the pre-trained models on Github at https://github.com/previtus [exact link TBR].